# IllustrisTNG in the HSC-SSP: No Shortage of Thin Disk Galaxies in TNG50

### v1

@@ Abstract
We perform a thorough analysis of the projected shapes of nearby galaxies in both observations and cosmological simulations. We implement a forward-modeling approach to overcome the limitations in previous studies which hinder accurate comparisons between observations and simulations. We measure axis ratios of $z=0$ (snapshot 99) TNG50 galaxies from their synthetic Hyper Suprime-Cam Subaru Strategic Program (HSC-SSP) images and compare them with those obtained from real HSC-SSP images of a matched galaxy sample. Remarkably, the comparison shows excellent agreement between observations and the TNG50 simulation. Specifically, for galaxies with stellar masses $10\leq\log(M_{\star}/M_{\sun})\leq 11.5$, we find $\lesssim 0.1\sigma$ tensions between the observation and the simulation, whereas previous studies found $\gtrsim 10\sigma$ tensions due to lack of thin galaxies in simulations. We reveal that low-mass galaxies ($M_{\star}\lesssim 10^{9.5}$) in TNG50 are thicker than their observed counterparts in HSC-SSP and attribute this to the spurious dynamical heating effects that artificially puff up galaxies. We also find that despite the overall broad agreement, TNG50 galaxies are more concentrated than the HSC-SSP ones at the low- and high-mass end of the stellar mass range of $9.0\leq\log(M_{\star}/M_{\sun})\leq 11.2$ and are less concentrated at intermediate stellar masses. But we argue that the higher concentrations of the low-mass TNG50 galaxies are not likely the cause of their thicker/rounder appearances. Our study underscores the critical importance of conducting mock observations of simulations and applying consistent measurement methodologies to facilitate proper comparison with observations.
@@ 1 Introduction
Comparing observations with cosmological predictions is a critical test of galaxy formation and evolution theories and also helps the interpretation of the observations. In addition to reproducing the fundamental scaling relations of galaxies, examining the detailed galaxy morphologies adds to the comprehensive measure of the accuracy and success of the models. Weinzirl et al. (2009) found that bulges in semi-analytic $\Lambda$CDM models are more prominent than the observed ones. Moreover, other studies claimed that there were an excessive number of bulgeless galaxies and galaxies with pseudo bulges compared to what they expected from hierarchical clustering (Kormendy & Fisher, 2008; Kormendy et al., 2010). However, it is worth noting that statistics of bulgeless galaxies and pseudo bulges are sensitive to the measurement techniques and selection criteria (Gao et al., 2018, 2019, 2020, 2022). Recent developments in hydrodynamical cosmological simulations, such as Illustris (Vogelsberger et al., 2013; Torrey et al., 2014), IllustrisTNG (hereafter TNG; Weinberger et al., 2017; Pillepich et al., 2018), and EAGLE (Crain et al., 2015; Schaye et al., 2015), have addressed the issue of angular momentum loss and produced more realistic looking galaxies (e.g., Vogelsberger et al., 2014). Still, tensions remain between the theories and observations, probably due to limitations in numerical resolution and physical models. Studies of both Illustris and EAGLE reported a significant deficit of bulge-dominated galaxies in the simulations (Bottrell et al., 2017a, b; de Graaff et al., 2022). On the contrary, Haslbauer et al. (2022; hereafter H22) reported a lack of thin galaxies in Illustris, TNG, and EAGLE, despite that their median axis ratios as a function of stellar mass agree well with observations (Rodriguez-Gomez et al., 2019). Adding to this, van de Sande et al. (2019) showed a lack of highly flattened galaxies in various simulations including EAGLE, HYDRANGEA (Bahé et al., 2017), and HORIZON-AGN (Dubois et al., 2014), except for Magneticum Pathfinder (Hirschmann et al., 2014; Teklu et al., 2015). The consistencies and conflicts demand more rigorous and self-consistent comparisons between observations and simulations to better understand the origins of such discrepancies.
@@
Some studies suffer from inconsistent treatment of the observational data and the simulations. Simulations offer 6D phase space information of the basic elements (e.g., stellar/dark matter particles and gas cells), while in observations only projected distribution and line-of-sight velocities of baryons are directly available. Because of the difference in availability of information and formats of data, authors often apply different sample selection criteria and measurement techniques to both sides, which helps to maximize the outcome from both but might have resulted in inaccurate conclusions. For example, H22 applied different methods to measure galaxy shapes from simulations and observations, which may have significant impacts on the fairness of the comparison. Meanwhile, significant effort has been devoted to ensuring the measurement methods are consistent; however, the consistency is often compromised by lack of realism or unavailability of mock observation of the simulations (e.g., van de Sande et al., 2019). Moreover, de Graaff et al. (2022) showed that using synthetic data of full realism helps alleviate the tension between the simulated and observed galaxy mass–size relation. We are therefore motivated to carry out a self-consistent comparison of galaxy shapes using synthetic images of simulated galaxies with full realism and aim to permanently resolve the controversy.
@@
Synthetic data have long been used to bridge the gap between simulations and observations for the purpose of comparison (e.g., Bottrell et al., 2017a, b; Rodriguez-Gomez et al., 2019; Eisert et al., 2024). In particular, radiative transfer post-processing of randomly oriented simulated galaxies to create mock images offers several advantages over previous studies: (1) we do not need to make assumptions of the 3D structure of the simulated galaxies for projection, (2) they automatically resolve the inconsistencies concerning mass-to-light ratios and instrumental effects, and (3) they allow application of the same measurement techniques to both sides. We will use a set of synthetic Hyper Suprime-Cam Subaru Strategic Program Survey (HSC-SSP; Aihara et al., 2018) images of TNG50 galaxies with $\log(M_{\star}/M_{\sun})\geq 9$ in the redshift range $0\leq z\leq 0.7$ created by Bottrell et al. (2024). Data and sample selection are described in Section 2. To mitigate systematics, we apply the same method to measure axis ratios of galaxies from real HSC-SSP images and the synthetic HSC-SSP images of TNG50, as outlined in Section 3. Results and discussions are given in Sections 4 and 5. We summarize the paper in Section 6. We adopt a cosmology based on Planck 2015 (Planck Collaboration et al., 2016) with $H_{0}=100h$ km s-1 Mpc-1 with $h=0.6774$, $\Omega_{m}=0.308$, and $\Omega_{\Lambda}=0.692$.
@@ 2.1 The Hyper Suprime-Cam Subaru Strategic Program Survey
The HSC-SSP (Aihara et al., 2018) is an optical imaging survey that uses the Hyper Suprime-Camera (Miyazaki et al., 2012, 2018) of the $8.2$-meter Subaru Telescope. HSC-SSP has released three public data releases (PDRs) (Aihara et al., 2018, 2019, 2022) and covers about $1000\,\mathrm{deg^{2}}$ in $grizy$ bands in the WIDE layer. One can reliably extract the surface brightness profiles of galaxies down to $28.5\,\mathrm{mag\,arcsec^{-2}}$ in $r$-band and $i$-band using PDR2 WIDE images (Huang et al., 2018; Li et al., 2022). The best spatial resolution is in the $i$-band, with a median seeing of $0\farcs 61$ for the PDR3 WIDE layer. The HSC has a pixel scale of $0\farcs 168$.
@@ 3.2 Measurements of Global Axis Ratios
The initialization is done in two steps. First, Fast Fourier Transform (FFT) is performed over fluxes along circular paths concentric with the galaxy and the phases are determined from the second Fourier coefficient $\mathcal{F}_{2}$. The global PA is the average direction of the phases for the five outermost circular paths whose average fluxes are still above the background level. Second, with the global PA being fixed, the global axis ratio is obtained by minimizing $\left|\mathcal{F}_{2}/\left(\tilde{\mu}+\sigma_{b}\right)\right|$, where $\tilde{\mu}$ is the median flux, $\sigma_{b}$ is the background noise, and $\tilde{\mu}+\sigma_{b}$ is approximately equal to $\mathcal{F}_{0}$.
The checking fit step performs four checks on the quality of the fits by examining the flux distribution along the best-fit isophotes: (1) and (2) the magnitudes of the first and the second Fourier mode, respectively; (3) the variability of fluxes; and (4) the variability of fluxes compared with that along the initialized isophotes. The first check failing due to significant power in the first Fourier mode often indicates a failed center determination or lopsidedness in galaxy flux distribution, and the second suggests that the optimization was not achieved or presence of elongated structures (e.g., bars). The third check compares the interquartile range of the fluxes to their median and fails when a minimum fraction of the isophotes have such high variabilities. The fourth check is similar to the third, only that the variability is contrasted against that obtained from the initialized isophotes. The thresholds for triggering the failures were chosen by Stone et al. (2021) to catch visually identified cases.
@@ 3.3 Measurements of Structural Parameters
We adopt the curve-of-growth (CoG) method to derive the concentration index for both the HSC-SSP galaxies and the TNG50 galaxies. We use AutoProf to extract 1D, azimuthally averaged surface brightness profiles of the galaxies. Subsequently, based on the growth curve of their flux truncated at the limiting surface brightness, we derive $r_{20}$ and $r_{80}$, where $r_{20}$ ($r_{80}$) represents the radius within which the integrated flux constitutes $20\%$ ($80\%$) of the total galaxy flux. We define the concentration index as $C_{82}=5\times\log{(r_{80}/r_{20})}$. We exclude galaxies with $r_{20}<1.5\times{\rm FWHM}$ to ensure sufficient spatial resolution for accurate measurements. The index characterizes the degree of flux concentration of a galaxy and is often used to separate bulge-dominated and disk-dominated galaxies (e.g., Conselice, 2003, 2014).
@@ 4.1 Comparison of Overall Axis Ratio Distributions between HSC-SSP and TNG50
where $N_{\mathrm{sim},i}$ denotes the number of simulated galaxies in the $i$-th axis ratio bin and $N^{\prime}_{\mathrm{obs},i}$ the weighted number of observed galaxies therein; $\sigma_{i}$ are their respective Poisson uncertainties. Following the modified Neyman’s $\chi^{2}$, when counts of galaxies are zero, a minimum counts of one is enforced to avoid the situation of $\sigma_{i}=0$. $N_{\mathrm{obs},i,j}$ is the number of observed galaxies in the $i$-th axis ratio bin and $j$-th mass bin. We will clarify the definition of their weights $w_{\mathrm{obs},j}$ when describing the details of the stellar-mass weighting scheme. $N_{{\rm bins},q}=20$ is the number of axis ratio bins with a fixed width of 0.05 and $N_{{\rm bins},M_{\star}}=10$ is the number of stellar mass bins with a fixed width of 0.15 dex.
@@
The different stellar mass distributions of the observed and the simulated galaxies should be taken into account when comparing their axis ratios (see Figure 1), as galaxy shape is known to depend on stellar mass. We follow the same stellar mass-weighting method in H22 to correct the axis ratio distribution of the observed galaxies before comparing it with that of the simulated galaxies. We calculate a weight for each observed galaxy as the number ratio of simulated galaxies to observed galaxies in the corresponding stellar mass bin. Specifically, the weight of an observed galaxy in the $j$-th stellar mass bin is defined as $w_{\mathrm{\rm obs},j}=N_{\mathrm{sim},j}/N_{\mathrm{obs},j}$. In the $i$-th axis ratio bin, $N^{\prime}_{\mathrm{obs},i}=\sum_{j}{w_{\mathrm{obs},j}N_{\mathrm{obs},i,j}}$ represents the total weight of galaxies therein, a summation of their $w_{{\rm obs},j}$. So, $N^{\prime}_{\mathrm{obs},i}$ is the adjusted number of the observed HSC-SSP galaxies in the axis ratio bin after the stellar-mass weighting, as if the observed galaxies obey the same stellar mass distribution of the TNG50 galaxies. The total weight of all the observed galaxies by definition equals the total number of simulated galaxies.
@@
This work HSC-SSP PDR3, $g$-band TNG50, synthetic $g$-band $4.44$ 0.00σ 1.00
This work HSC-SSP PDR3, $i$-band TNG50, synthetic $i$-band $8.77$ 0.03σ 0.98
This work HSC-SSP PDR3, $y$-band TNG50, synthetic $y$-band $11.83$ 0.13σ 0.89
@@ 4.2 Comparison of Axis Ratio Distributions between HSC-SSP and TNG50 at Fixed Stellar Masses
Building on the remarkable consistency in the overall axis ratio distributions between the observations and TNG50, this section will delve deeper into exploring the dependence of axis ratios on stellar mass over a broader stellar mass range. The $q\textendash\log M_{\star}$ diagram shown in Figure 4 juxtaposes the galaxies in HSC-SSP and TNG50 with stellar masses of $9.0\leq\log(M_{\star}/M_{\sun})\leq 11.2$. We truncate the sample at $\log(M_{\star}/M_{\sun})=11.2$ to ensure enough galaxies at the highest stellar mass for decent sampling of their axis ratio distribution. To simulate the uncertainties in real-world stellar mass estimates, we also randomly perturb the stellar masses of the TNG50 galaxies following Rodriguez-Gomez et al. (2019) and show the results in the rightmost column of Figure 4. To illustrate the dependence of the axis ratio distributions on stellar mass, we show the 1st, 16th, 50th, 84th, and 99th percentiles of the axis ratios in each stellar mass bin of $0.25$ dex width. The percentiles of the HSC-SSP and TNG50 galaxies show broad agreement with each other. However, we note that there are noticeable differences at lower stellar masses ($\log(M_{\star}/M_{\sun})\lesssim 9.5$) and in redder bands ($y$ band). We perform similar comparisons as those in Section 4.1 but in much smaller stellar mass bins and find that the difference between the axis ratio distributions of the HSC-SSP galaxies and the TNG50 galaxies are generally small ($\leq 2\sigma$) except for the $y$-band axis ratios in the stellar mass range $\log(M_{\star}/M_{\sun})<9.5$, which surges to $>8\sigma$ (Table 2). Moreover, the statistical significance of the differences at low stellar mass increases monotonically towards longer wavelengths that better probe stellar structures of galaxies, although the increase from $g$ to $i$ band is marginal compared to that from $i$ to $y$ band. The differences in the $y$-band axis ratios are still significant even after smoothing the trend with perturbing the TNG50 stellar masses.

### v2

@@ Abstract
We perform a thorough analysis of the projected shapes of nearby galaxies in both observations and cosmological simulations. We implement a forward-modeling approach to overcome the limitations in previous studies, which hinder accurate comparisons between observations and simulations. We measure axis ratios of $z=0$ (snapshot 99) TNG50 galaxies from their synthetic Hyper Suprime-Cam Subaru Strategic Program (HSC-SSP) images and compare them with those obtained from real HSC-SSP images of a matched galaxy sample. Remarkably, the comparison shows excellent agreement between the observations and the TNG50 simulation, challenging previous claims that $\Lambda$CDM models underproduced the abundance of thin galaxies. Specifically, for galaxies with stellar masses $10\leq\log(M_{\star}/M_{\sun})\leq 11.5$, we find $\lesssim 0.1\sigma$ tensions between the observations and the simulation, a stark contrast to the previously reported $\gtrsim 10\sigma$ tensions. We reveal that low-mass galaxies ($M_{\star}\lesssim 10^{9.5}\,M_{\odot}$) in TNG50 are thicker than their observed counterparts in HSC-SSP and attribute this to the spurious dynamical heating effects that artificially puff up galaxies. We also find that, despite the overall broad agreement, TNG50 galaxies are more concentrated than the HSC-SSP ones at the low- and high-mass end of the stellar mass range of $9.0\leq\log(M_{\star}/M_{\sun})\leq 11.2$ and are less concentrated at intermediate stellar masses. But we argue that the higher concentrations of the low-mass TNG50 galaxies are not likely the cause of their thicker/rounder appearances. Our study underscores the critical importance of conducting mock observations of simulations and applying consistent measurement methodologies to facilitate proper comparison with observations.
@@ 1 Introduction
Comparing observations with cosmological predictions is a critical test of galaxy formation and evolution theories and also helps the interpretation of the observations. In addition to reproducing the fundamental scaling relations of galaxies, examining the detailed galaxy morphologies adds to the comprehensive measure of the accuracy and success of the models. Weinzirl et al. (2009) found that bulges in semi-analytic $\Lambda$CDM models are more prominent than the observed ones. Moreover, other studies claimed that there were an excessive number of bulgeless galaxies and galaxies with pseudo bulges compared to what they expected from hierarchical clustering (Kormendy & Fisher, 2008; Kormendy et al., 2010). However, it is worth noting that statistics of bulgeless galaxies and pseudo bulges are sensitive to the measurement techniques and selection criteria (Gao & Ho, 2017; Gao et al., 2018, 2019, 2020, 2022). Recent developments in hydrodynamical cosmological simulations, such as Illustris (Vogelsberger et al., 2013; Torrey et al., 2014), IllustrisTNG (hereafter TNG; Weinberger et al., 2017; Pillepich et al., 2018), and EAGLE (Crain et al., 2015; Schaye et al., 2015), have addressed the issue of angular momentum loss and produced more realistic looking galaxies (e.g., Vogelsberger et al., 2014). Still, tensions remain between the theories and observations, probably due to limitations in numerical resolution and physical models. Studies of both Illustris and EAGLE reported a significant deficit of bulge-dominated galaxies in the simulations (Bottrell et al., 2017a, b; de Graaff et al., 2022). On the contrary, Haslbauer et al. (2022; hereafter H22) reported a lack of thin galaxies in Illustris, TNG, and EAGLE, despite that their median axis ratios as a function of stellar mass agree well with observations (Rodriguez-Gomez et al., 2019). Adding to this, van de Sande et al. (2019) showed a lack of highly flattened galaxies in various simulations including EAGLE, HYDRANGEA (Bahé et al., 2017), and HORIZON-AGN (Dubois et al., 2014), except for Magneticum Pathfinder (Hirschmann et al., 2014; Teklu et al., 2015). The consistencies and conflicts demand more rigorous and self-consistent comparisons between observations and simulations to better understand the origins of such discrepancies.
@@
Some studies suffer from inconsistent treatment of the observational data and the simulations. Simulations offer 6D phase space information of the basic elements (e.g., stellar/dark matter particles and gas cells), while in observations only projected distribution and line-of-sight velocities of baryons are directly available. Because of the difference in availability of information and formats of data, authors often apply different sample selection criteria and measurement techniques to both sides, which helps to maximize the outcome from both but might have resulted in inaccurate conclusions. For example, H22 applied different methods to measure galaxy shapes from simulations and observations, which may have significant impacts on the fairness of the comparison. Meanwhile, significant effort has been devoted to ensuring the measurement methods are consistent; however, the consistency is often compromised by lack of realistic mock observation of the simulations (e.g., van de Sande et al., 2019). Moreover, de Graaff et al. (2022) showed that using synthetic data of full realism helps alleviate the tension between the simulated and observed galaxy mass–size relation. We are therefore motivated to carry out a self-consistent comparison of galaxy shapes using synthetic images of simulated galaxies with full realism and aim to permanently resolve the controversy.
@@
Synthetic data have long been used to bridge the gap between simulations and observations for the purpose of comparison (e.g., Bottrell et al., 2017a, b; Rodriguez-Gomez et al., 2019; Eisert et al., 2024). In particular, radiative transfer post-processing of randomly oriented simulated galaxies to create mock images offers several advantages over previous studies: (1) we do not need to make assumptions of the 3D structure of the simulated galaxies for projection, (2) they automatically resolve the inconsistencies concerning mass-to-light ratios and instrumental effects, and (3) they allow application of the same measurement techniques to both sides. We will use a set of synthetic Hyper Suprime-Cam Subaru Strategic Program (HSC-SSP; Aihara et al., 2018) images of TNG50 galaxies with $\log(M_{\star}/M_{\sun})\geq 9$ in the redshift range $0\leq z\leq 0.7$ created by Bottrell et al. (2024). Data and sample selection are described in Section 2. To mitigate systematics, we apply the same method to measure axis ratios of galaxies from real HSC-SSP images and the synthetic HSC-SSP images of TNG50, as outlined in Section 3. Results and discussions are given in Sections 4 and 5. We summarize the paper in Section 6. We adopt a cosmology based on Planck 2015 (Planck Collaboration et al., 2016) with $H_{0}=100h$ km s-1 Mpc-1 with $h=0.6774$, $\Omega_{m}=0.308$, and $\Omega_{\Lambda}=0.692$.
@@ 2.1 The Hyper Suprime-Cam Subaru Strategic Program Survey
The HSC-SSP (Aihara et al., 2018) is an optical imaging survey that uses the HSC (Miyazaki et al., 2012, 2018) of the $8.2$-meter Subaru Telescope. HSC-SSP has released three public data releases (PDRs; Aihara et al., 2018, 2019, 2022) and covers about $1000\,\mathrm{deg^{2}}$ in $grizy$ bands in the WIDE layer. One can reliably extract the surface brightness profiles of galaxies down to $28.5\,\mathrm{mag\,arcsec^{-2}}$ in $r$ band and $i$ band using PDR2 WIDE images (Huang et al., 2018; Li et al., 2022). The best spatial resolution is in the $i$-band, with a median seeing of $0\farcs 61$ for the PDR3 WIDE layer. The HSC has a pixel scale of $0\farcs 168$.
@@ 3.2 Measurements of Global Axis Ratios
The initialization is done in two steps. First, fast Fourier transform (FFT) is performed over fluxes along circular paths concentric with the galaxy, and the phases are determined from the second Fourier coefficient $\mathcal{F}_{2}$. The global PA is the average direction of the phases for the five outermost circular paths whose average fluxes are still above the background level. Second, with the global PA being fixed, the global axis ratio is obtained by minimizing $\left|\mathcal{F}_{2}/\left(\tilde{\mu}+\sigma_{b}\right)\right|$, where $\tilde{\mu}$ is the median flux, $\sigma_{b}$ is the background noise, and $\tilde{\mu}+\sigma_{b}$ is approximately equal to $\mathcal{F}_{0}$.
The checking fit step performs four checks on the quality of the fits by examining the flux distribution along the best-fit isophotes: (1) and (2) the magnitudes of the first and the second Fourier mode, respectively; (3) the variability of fluxes; and (4) the variability of fluxes compared with that along the initialized isophotes. The first check failing due to significant power in the first Fourier mode often indicates a failed center determination or lopsidedness in galaxy flux distribution, and the second suggests that the optimization was not achieved or the presence of elongated structures (e.g., bars). The third check compares the interquartile range of the fluxes to their median and fails when a minimum fraction of the isophotes have such high variabilities. The fourth check is similar to the third, only that the variability is contrasted against that obtained from the initialized isophotes. The thresholds for triggering the failures were chosen by Stone et al. (2021) to catch visually identified cases.
@@ 3.3 Measurements of Structural Parameters
We adopt the curve-of-growth method to derive the concentration index for both the HSC-SSP galaxies and the TNG50 galaxies. We use AutoProf to extract 1D, azimuthally averaged surface brightness profiles of the galaxies. Subsequently, based on the growth curve of their flux truncated at the limiting surface brightness, we derive $r_{20}$ and $r_{80}$, where $r_{20}$ ($r_{80}$) represents the radius within which the integrated flux constitutes $20\%$ ($80\%$) of the total galaxy flux. We define the concentration index as $C_{82}=5\times\log{(r_{80}/r_{20})}$. We exclude galaxies with $r_{20}<1.5\times{\rm FWHM}$ to ensure sufficient spatial resolution for accurate measurements. The index characterizes the degree of flux concentration of a galaxy and is often used to separate bulge-dominated and disk-dominated galaxies (e.g., Conselice, 2003, 2014).
@@ 4.1 Comparison of Overall Axis Ratio Distributions between HSC-SSP and TNG50
where $N_{\mathrm{sim},i}$ denotes the number of simulated galaxies in the $i$th axis ratio bin and $N^{\prime}_{\mathrm{obs},i}$ the weighted number of observed galaxies therein; $\sigma_{i}$ are their respective Poisson uncertainties. Following the modified Neyman’s $\chi^{2}$, when counts of galaxies are zero, a minimum counts of one is enforced to avoid the situation of $\sigma_{i}=0$. $N_{\mathrm{obs},i,j}$ is the number of observed galaxies in both the $i$th axis ratio bin and the $j$th stellar mass bin. We will clarify the definition of their weights $w_{\mathrm{obs},j}$ when describing the details of the stellar-mass weighting scheme. $N_{{\rm bins},q}=20$ is the number of axis ratio bins with a fixed width of 0.05 and $N_{{\rm bins},M_{\star}}=10$ is the number of stellar mass bins with a fixed width of 0.15 dex.
@@
The different stellar mass distributions of the observed and the simulated galaxies should be taken into account when comparing their axis ratios (see Figure 1), as galaxy shape is known to depend on stellar mass. We follow the same stellar mass-weighting method in H22 to correct the axis ratio distribution of the observed galaxies before comparing it with that of the simulated galaxies. We calculate a weight for each observed galaxy as the number ratio of simulated galaxies to observed galaxies in the corresponding stellar mass bin. Specifically, the weight of an observed galaxy in the $j$th stellar mass bin is defined as $w_{\mathrm{\rm obs},j}=N_{\mathrm{sim},j}/N_{\mathrm{obs},j}$. In the $i$th axis ratio bin, $N^{\prime}_{\mathrm{obs},i}=\sum_{j}{w_{\mathrm{obs},j}N_{\mathrm{obs},i,j}}$ represents the total weight of galaxies therein, a summation of their $w_{{\rm obs},j}$. So, $N^{\prime}_{\mathrm{obs},i}$ is the adjusted number of the observed HSC-SSP galaxies in the axis ratio bin after the stellar-mass weighting, as if the observed galaxies obey the same stellar mass distribution of the TNG50 galaxies. The total weight of all the observed galaxies by definition equals the total number of simulated galaxies.
@@
This work HSC-SSP PDR3, $g$ band TNG50, synthetic $g$ band $4.44$ 0.00σ 1.00
This work HSC-SSP PDR3, $i$ band TNG50, synthetic $i$ band $8.77$ 0.03σ 0.98
This work HSC-SSP PDR3, $y$ band TNG50, synthetic $y$ band $11.8$ 0.13σ 0.89
@@ 4.2 Comparison of Axis Ratio Distributions between HSC-SSP and TNG50 at Fixed Stellar Masses
Building on the remarkable consistency in the overall axis ratio distributions between the observations and TNG50, this section will delve deeper into exploring the dependence of axis ratios on stellar mass over a broader stellar mass range. The $q\textendash\log M_{\star}$ diagram shown in Figure 4 juxtaposes the galaxies in HSC-SSP and TNG50 with stellar masses of $9.0\leq\log(M_{\star}/M_{\sun})\leq 11.2$. We truncate the sample at $\log(M_{\star}/M_{\sun})=11.2$ to ensure enough galaxies at the highest stellar mass for decent sampling of their axis ratio distribution. To simulate the uncertainties in real-world stellar mass estimates, we also randomly perturb the stellar masses of the TNG50 galaxies as suggested by Rodriguez-Gomez et al. (2019) and show the results in the rightmost column of Figure 4. To illustrate the dependence of the axis ratio distributions on stellar mass, we show the 1st, 16th, 50th, 84th, and 99th percentiles of the axis ratios in each stellar mass bin of $0.25$ dex width. The percentiles of the HSC-SSP and TNG50 galaxies show broad agreement with each other. However, we note that there are noticeable differences at lower stellar masses ($\log(M_{\star}/M_{\sun})\lesssim 9.5$) and in redder bands ($y$ band). We perform similar comparisons as those in Section 4.1 but in much smaller stellar mass bins and find that the difference between the axis ratio distributions of the HSC-SSP galaxies and the TNG50 galaxies are generally small ($\leq 2\sigma$) except for the $y$-band axis ratios in the stellar mass range $\log(M_{\star}/M_{\sun})<9.5$, which surges to $>8\sigma$ (Table 2). Moreover, the statistical significance of the differences at low stellar mass increases monotonically toward longer wavelengths that better probe stellar structures of galaxies, although the increase from $g$ to $i$ band is marginal compared to that from $i$ to $y$ band. The differences in the $y$-band axis ratios are still significant even after smoothing the trend with perturbing the TNG50 stellar masses.